OpenStudy (anonymous):

Prove that if n is an integer greater than 2, then n^3 - 1 is composite.

OpenStudy (mani_jha):

n^3-1=(n-1){1+n(n+1)} If n is odd, n-1 is even and n+1 is even. So, n(n+1) is also even. Thus, n(n+1)+1 is odd. So, multiplication of even no (n-1) and odd n(n+1)+1 gives an even number, which is composite. If n is even, n-1 is odd and n+1 is odd. So, n(n+1) is even. Thus, n(n+1)+1 is odd. So, multiplication of odd no (n-1) and odd n(n+1)+1 gives an odd number, which is composite.

OpenStudy (anonymous):

good job mani

OpenStudy (anonymous):

excellent explanation

OpenStudy (anonymous):

Thank you! :)

OpenStudy (mani_jha):

Thanks guys

OpenStudy (anonymous):

If \(n>2\) then \(n-1\geq 2\) which means that \[n^3-1=(n-1)(n^2+n+1)\] must be composite.